# The Seifert-van Kampen Theorem via Computational Paths: A Formalized Approach to Computing Fundamental Groups

### v2

@@ 3 Quot (AmalgEquiv i1 i2)
where $\mathrm{AmalgEquiv}$ is the equivalence closure of $\mathrm{AmalgRelation}$.
Remark 4.6 (Structure of AmalgEquiv).
The relation $\mathrm{AmalgEquiv}$ includes:
  1. (1)
    Standard word-reduction equivalences for the free product $G_{1}\ast G_{2}$: combining adjacent letters from the same factor, and removing identity elements.
  2. (2)
    The amalgamation relations identifying $i_{1}(h)$ with $i_{2}(h)$ for all $h:H$.
  3. (3)
    Reflexive, symmetric, and transitive closure.
In the Lean formalization, we build the free product quotient first, then extend with amalgamation. The result is the pushout in the category of groups.
Lemma 4.7 (Amalgamation Property).
For any $h:H$:
$\mathrm{inl}(i_{1}(h))=\mathrm{inr}(i_{2}(h))\quad\text{in }G_{1}\ast_{H}G_{2}$
Proof.
By definition, the singleton words $\mathrm{singleLeft}(i_{1}(h))$ and $\mathrm{singleRight}(i_{2}(h))$ are related by $\mathrm{AmalgRelation}$ (with empty prefix and suffix). The quotient $\mathrm{Quot.sound}$ then yields the equality. ∎
@@ 5. The Seifert-van Kampen Theorem
With pushouts (Section 3) and amalgamated free products (Section 4) in hand, we can now state and prove the main theorem.
@@ 5.1. Statement
Theorem 5.1 (Seifert-van Kampen).
@@
where the amalgamation is via the induced maps $f_{*}:\pi_{1}(C,c_{0})\to\pi_{1}(A,f(c_{0}))$ and $g_{*}:\pi_{1}(C,c_{0})\to\pi_{1}(B,g(c_{0}))$.
Proof sketch.
We construct:
  1. (1)
    $\mathrm{decode}$: Maps words to loops by interpreting left letters as loops in $\mathrm{inl}(A)$, and right letters as loops in $\mathrm{inr}(B)$ conjugated by $\mathrm{glue}(c_{0})$.
  2. (2)
    $\mathrm{encode}$: Defined via transport of a code family along paths (axiomatized).
The key lemmas are:
  • Decode respects amalgamation (Lemma 5.3): elements from $\pi_{1}(C)$ decode identically whether placed in the left or right factor.
  • Round-trip properties: $\mathrm{decode}\circ\mathrm{encode}=\mathrm{id}$ and $\mathrm{encode}\circ\mathrm{decode}\sim_{\mathrm{amalg}}\mathrm{id}$.
The first round-trip follows by path induction on pushout paths; the second by word induction. See §5.2.35.2.4 for details. ∎
@@ 5.2. The Encode-Decode Method
The proof uses the encode-decode method:
  1. (1)
    Decode: Define a map from the amalgamated free product to $\pi_{1}(\mathrm{Pushout})$.
  2. (2)
    Encode: Define a map from $\pi_{1}(\mathrm{Pushout})$ to the amalgamated free product.
  3. (3)
    Round-trip: Show both compositions are identity.
@@ 2 FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
6 let alphaInPushout := Quot.lift
7 (fun p => Quot.mk RwEq (Pushout.inlPath p))
8 (fun _ _ hp => Quot.sound (rweq_congrArg_of_rweq Pushout.inl hp))
9 alpha
10 piOneMul alphaInPushout (pushoutDecode c0 rest)
@@
12 let betaInPushout := Quot.lift
13 (fun q => Quot.mk RwEq (Path.trans
14 (Pushout.glue c0)
15 (Path.trans (Pushout.inrPath q)
16 (Path.symm (Pushout.glue c0)))))
17 (fun _ _ hq => Quot.sound (...))
18 beta
19 piOneMul betaInPushout (pushoutDecode c0 rest)
The key insight for elements from $\pi_{1}(B)$: since the basepoint is in $\mathrm{inl}(A)$, we must conjugate by the glue path to create a loop:
@@
$\mathrm{inl}(f(c_{0}))$$\mathrm{inr}(g(c_{0}))$$\mathrm{inl}(f(c_{0}))$$\mathrm{inr}(g(c_{0}))$$\scriptstyle{\mathrm{glue}(c_{0})}$$\scriptstyle{\beta_{\mathrm{decoded}}}$$\scriptstyle{\mathrm{inr}_{*}(\beta)}$$\scriptstyle{\mathrm{glue}(c_{0})^{-1}}$
@@ 5.2.2. Decode Respects Amalgamation
We need to show that for a loop $p$ in $C$ at $c_{0}$:
$\mathrm{inl}_{*}(f_{*}(p))\sim\mathrm{glue}(c_{0})\cdot\mathrm{inr}_{*}(g_{*}(p))\cdot\mathrm{glue}(c_{0})^{-1}$
This is exactly Corollary 3.4 applied to the loop $p$ (where $c_{1}=c_{2}=c_{0}$):
1theorem glue_natural_loop_rweq (c0 : C) (p : LoopSpace C c0) :
2 RwEq (inlPath (Path.congrArg f p))
3 (Path.trans (glue c0)
4 (Path.trans (inrPath (Path.congrArg g p))
5 (Path.symm (glue c0)))) :=
6 glue_natural_rweq p
@@ 5.2.3. The Encode Map
The encode map uses the “code family” approach, a technique central to HoTT calculations of fundamental groups [10, 17]. The idea is to define a type family $\mathrm{Code}:\mathrm{Pushout}\to\mathrm{Type}$ that, at each point, represents the words that “should decode to a path ending at that point.”
  • At $\mathrm{inl}(a)$: $\mathrm{Code}(\mathrm{inl}(a))$ represents words decoding to paths ending at $\mathrm{inl}(a)$
  • At $\mathrm{inr}(b)$: $\mathrm{Code}(\mathrm{inr}(b))$ represents words decoding to paths ending at $\mathrm{inr}(b)$
  • Transport along $\mathrm{glue}(c)$: adjusts words by prepending/removing letters for the glue path
The encode function is then:
$\mathrm{encode}(p):=\mathrm{transport}(\mathrm{Code},p,\mathrm{nil})$
Why encode is axiomatized.
Fully constructing the code family requires dependent elimination over the pushout HIT, which in Lean 4 (without native HIT support or cubical features) would require extensive additional infrastructure. Since our focus is on the computational paths framework, we axiomatize encode’s existence and key properties:
1axiom pushoutEncodeAxiom (A B C : Type u) (f : C -> A) (g : C -> B) (c0 : C) :
2 LoopSpace (Pushout A B C f g) (Pushout.inl (f c0)) ->
3 FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
4
5axiom pushoutEncodeAxiom_respects_rweq (A B C : Type u)
6 (f : C -> A) (g : C -> B) (c0 : C)
7 {p q : LoopSpace (Pushout A B C f g) (Pushout.inl (f c0))}
8 (h : RwEq p q) :
9 pushoutEncodeAxiom A B C f g c0 p = pushoutEncodeAxiom A B C f g c0 q
The second axiom ensures encode respects $\sim$, essential for lifting to the $\pi_{1}$ quotient. In cubical type theory, these would be provable; here, they capture the interface needed for SVK.
@@ 5.2.4. Round-Trip Properties
Lemma 5.4 (Decode-Encode Round-Trip).
For any loop $p$ in the pushout at $\mathrm{inl}(f(c_{0}))$:
$\mathrm{decode}(\mathrm{encode}(p))=[p]$
Lemma 5.5 (Encode-Decode Round-Trip).
For any word $w$:
$\mathrm{encode}(\mathrm{decode}(w))\sim_{\mathrm{amalg}}w$
where $\sim_{\mathrm{amalg}}$ is amalgamation equivalence.
5.3. The Equivalence
The equivalence is defined in Lean as:
1noncomputable def seifertVanKampenEquiv :
2 SimpleEquiv
3 (pi_1(Pushout A B C f g, Pushout.inl (f c0)))
4 (AmalgamatedFreeProduct
5 (pi_1(A, f c0))
6 (pi_1(B, g c0))
7 (pi_1(C, c0))
8 (piOneFmap c0)
9 (piOneGmap c0)) where
10 toFun := pushoutEncodeAmalg c0
11 invFun := pushoutDecodeAmalg c0
12 left_inv := by
13 intro alpha
14 induction alpha using Quot.ind with
15 | _ p => exact pushoutDecodeEncodeAxiom A B C f g c0 p
16 right_inv := by
17 intro x
18 induction x using Quot.ind with
19 | _ w =>
20 apply Quot.sound
21 exact pushoutEncodeDecodeAxiom A B C f g c0 w
@@ 6.1. The Wedge Sum: Free Product of Fundamental Groups
When the gluing space $C$ is a point, $\pi_{1}(C)$ is trivial and the amalgamated free product reduces to a true free product.
@@
The wedge $A\vee B=\mathrm{Pushout}(A,B,\mathbf{1},\lambda\_.a_{0},\lambda\_.b_{0})$. By SVK:
$\pi_{1}(A\vee B)\simeq\pi_{1}(A)\ast_{\pi_{1}(\mathbf{1})}\pi_{1}(B)$
Since $\pi_{1}(\mathbf{1})\cong 1$ (the trivial group), the amalgamation imposes no relations, yielding the free product. ∎
@@
1noncomputable def wedgeFundamentalGroupEquiv :
2 SimpleEquiv
3 (pi_1(Wedge A B a0 b0, Wedge.basepoint))
4 (FreeProductWord (pi_1(A, a0)) (pi_1(B, b0))) where
5 toFun := wedgeEncodeQuot a0 b0
6 invFun := wedgeFreeProductDecode a0 b0
7 left_inv := ...
8 right_inv := ...
@@ 6.2. The Figure-Eight Space
Proof.
The figure-eight is the wedge of two circles: $\mathrm{FigureEight}=S^{1}\vee S^{1}$.
By Theorem 6.1:
$\pi_{1}(S^{1}\vee S^{1})\simeq\pi_{1}(S^{1})\ast\pi_{1}(S^{1})\simeq\mathbb{Z}\ast\mathbb{Z}$
using the fundamental theorem $\pi_{1}(S^{1})\simeq\mathbb{Z}$ [10]. ∎
In Lean:
1def FigureEight : Type u := Wedge Circle Circle circleBase circleBase
2
3noncomputable def figureEightPiOneEquivPiOneProduct :
4 SimpleEquiv FigureEightPiOne (FreeProductWord circlePiOne circlePiOne) :=
5 wedgeFundamentalGroupEquiv circleBase circleBase
6
7noncomputable def figureEightPiOneEquiv :
8 SimpleEquiv FigureEightPiOne (FreeProductWord Int Int) :=
9 SimpleEquiv.comp figureEightPiOneEquivPiOneProduct piOneProductEquivIntProduct
6.2.1. Non-Abelianness
The figure-eight has a non-abelian fundamental group, witnessed by the word structure:
@@ 6.2.1. Non-Abelianness
5 intro h
6 cases h -- consLeft and consRight are distinct constructors
This shows that $\mathrm{loop}_{A}\cdot\mathrm{loop}_{B}\neq\mathrm{loop}_{B}\cdot\mathrm{loop}_{A}$. The proof is by case analysis: FreeProductWord is an inductive type where consLeft and consRight are distinct constructors, so no sequence of reductions can make them equal.
Topologically, $\mathrm{wordAB}$ corresponds to first traversing the left circle, then the right; $\mathrm{wordBA}$ traverses them in the opposite order. In the figure-eight, these loops cannot be continuously deformed into each other—they are genuinely distinct homotopy classes.
6.3. The 2-Sphere: Trivial Fundamental Group
Theorem 6.3 (Fundamental Group of 2-Sphere).
$\pi_{1}(S^{2})\simeq 1$
@@ 6.3. The 2-Sphere: Trivial Fundamental Group
The 2-sphere is the suspension of the circle: $S^{2}=\Sigma(S^{1})$.
As a pushout: $\Sigma(S^{1})=\mathrm{Pushout}(\mathbf{1},\mathbf{1},S^{1},!,!)$ where !! is the unique map to the unit type.
By SVK:
$\pi_{1}(S^{2})\simeq\pi_{1}(\mathbf{1})\ast_{\pi_{1}(S^{1})}\pi_{1}(\mathbf{1})=1\ast_{\mathbb{Z}}1=1$
The key insight: when both factors are trivial, the amalgamated product is trivial regardless of the amalgamating group. ∎
Lemma 6.4 (Trivial Decode).
When $\pi_{1}(A)=\pi_{1}(B)=1$, for any word $w$:
$\mathrm{decode}(w)=[\rho]$

### v3

@@ 6 Quot (AmalgEquiv i1 i2)
4.3. Full Amalgamated Free Product
For a “canonical” target with fully reduced words, we define the full amalgamated free product that additionally reduces via free group laws:
1inductive FreeGroupStep : FreeProductWord G1 G2 -> FreeProductWord G1 G2 -> Prop
2 | cancel_left (x : G1) (pre suf : FreeProductWord G1 G2)
3 (h : x + (-x) = 0) :
4 FreeGroupStep (concat pre (consLeft x (consLeft (-x) suf)))
5 (concat pre suf)
6 | cancel_right (y : G2) (pre suf : FreeProductWord G1 G2)
7 (h : y + (-y) = 0) :
8 FreeGroupStep (concat pre (consRight y (consRight (-y) suf)))
9 (concat pre suf)
10
11def FullAmalgEquiv := EqvGen (fun w1 w2 =>
12 AmalgRelation i1 i2 w1 w2 \/ FreeGroupStep w1 w2)
13
14def FullAmalgamatedFreeProduct (G1 G2 H : Type u) (i1 : H -> G1) (i2 : H -> G2) :=
15 Quot (FullAmalgEquiv i1 i2)
Remark 4.5 (Standard vs Full Target).
The AmalgamatedFreeProduct is the standard target where words are only quotiented by amalgamation. The FullAmalgamatedFreeProduct additionally applies free group reductions. Both are valid SVK targets; the choice depends on application needs.
@@ 5.1. Statement
Theorem 5.1 (Seifert-van Kampen Schema).
@@
Dependency Summary for Theorem 5.1 Proved: The decode map $\mathrm{decode}:\mathrm{Word}\to\pi_{1}(\mathrm{Pushout})$ is constructed explicitly. Proved: Decode respects the amalgamation relation (Lemma 5.2), assuming glue naturality (HasGlueNaturalLoopRwEq). Proved: The quotient map $\mathrm{decode}_{\mathrm{Amalg}}:\mathrm{AmalgFP}\to\pi_{1}(\mathrm{Pushout})$ is well-defined. $\square$ Assumed (typeclass): The encode map exists (HasPushoutSVKEncodeQuot). $\square$ Assumed (typeclass): Decode-encode round-trip (HasPushoutSVKDecodeEncode). $\square$ Assumed (typeclass): Encode-decode round-trip (HasPushoutSVKEncodeDecode). The equivalence is therefore parametric in user-supplied encode/decode structure. Alternatively, the choice-based interface (HasPushoutSVKDecodeAmalgBijective) requires only a Prop-level bijectivity assumption without explicit encode.
5.2. Split Assumptions Architecture
The SVK proof is structured around split typeclass assumptions, allowing maximum flexibility:
1-- Just the encode map
2class HasPushoutSVKEncodeQuot (c0 : C) where
3 pushoutEncodeQuot : pi_1(Pushout A B C f g, inl (f c0)) ->
4 AmalgamatedFreeProduct (pi_1 A) (pi_1 B) (pi_1 C)
5
6-- Decode-encode round-trip (Prop)
7class HasPushoutSVKDecodeEncode (c0 : C) where
8 decode_encode : forall alpha, pushoutDecodeAmalg c0 (encode alpha) = alpha
9
10-- Encode-decode round-trip (up to AmalgEquiv)
11class HasPushoutSVKEncodeDecode (c0 : C) where
12 encode_decode : forall w, AmalgEquiv (encode (decode w)) w
13
14-- Encode-decode up to FullAmalgEquiv
15class HasPushoutSVKEncodeDecodeFull (c0 : C) where
16 encode_decode_full : forall w, FullAmalgEquiv (encode (decode w)) w
17
18-- Prop-level bijectivity (choice-based)
19class HasPushoutSVKDecodeAmalgBijective (c0 : C) where
20 decode_injective : Function.Injective (pushoutDecodeAmalg c0)
21 decode_surjective : Function.Surjective (pushoutDecodeAmalg c0)
5.3. Multiple Interfaces
The framework provides multiple equivalences depending on available assumptions:
1-- Standard SVK with explicit encode data
2noncomputable def seifertVanKampenEquiv
3 [HasPushoutSVKEncodeQuot c0] [HasGlueNaturalLoopRwEq c0]
4 [HasPushoutSVKDecodeEncode c0] [HasPushoutSVKEncodeDecode c0] :
5 SimpleEquiv (pi_1(Pushout A B C f g, inl (f c0)))
6 (AmalgamatedFreeProduct (pi_1 A) (pi_1 B) (pi_1 C))
7
8-- SVK with full target (free reduction)
9noncomputable def seifertVanKampenFullEquiv
10 [HasPushoutSVKEncodeQuot c0] [HasGlueNaturalLoopRwEq c0]
11 [HasPushoutSVKDecodeEncode c0] [HasPushoutSVKEncodeDecodeFull c0] :
12 SimpleEquiv (pi_1(Pushout A B C f g, inl (f c0)))
13 (FullAmalgamatedFreeProduct (pi_1 A) (pi_1 B) (pi_1 C))
14
15-- Choice-based SVK (no explicit encode required)
16noncomputable def seifertVanKampenEquiv_of_decodeAmalg_bijective
17 [HasPushoutSVKDecodeAmalgBijective c0] [HasGlueNaturalLoopRwEq c0] :
18 SimpleEquiv (pi_1(Pushout A B C f g, inl (f c0)))
19 (AmalgamatedFreeProduct (pi_1 A) (pi_1 B) (pi_1 C))
@@ 2 FreeProductWord (pi_1(A, f c0)) (pi_1(B, g c0))
6 piOneMul (liftLeft alpha) (pushoutDecode c0 rest)
@@
8 piOneMul (conjugateRight beta) (pushoutDecode c0 rest)
where conjugateRight conjugates by the glue path:
@@ 5.4.2. Decode Respects Amalgamation
This follows from glue naturality: $\mathrm{inl}_{*}(f_{*}(p))\sim\mathrm{glue}(c_{0})\cdot\mathrm{inr}_{*}(g_{*}(p))\cdot\mathrm{glue}(c_{0})^{-1}$. ∎
In Lean:
1theorem pushoutDecodeAmalg_respects_rel [HasGlueNaturalLoopRwEq c0]
2 (h : HasAmalgRelation i1 i2 w1 w2) :
3 pushoutDecodeAmalg c0 w1 = pushoutDecodeAmalg c0 w2 := by
4 induction h with
5 | amalgLeftToRight gamma pre suf =>
6 -- Use glue naturality to show the images are equal
7 simp only [pushoutDecodeAmalg]
8 apply congrArg
9 -- Key step: glue_natural_loop shows the conjugation
10 exact glue_natural_loop_pi1 gamma
@@ 5.4.3. The Encode Map
The encode map is axiomatized via the typeclass HasPushoutSVKEncodeQuot. In a cubical type theory, this would be constructed via the code family method.
@@ 5.4.4. Round-Trip Properties
The round-trip properties establish the equivalence:
  • $\mathrm{decode}(\mathrm{encode}(\alpha))=\alpha$ (via HasPushoutSVKDecodeEncode)
  • $\mathrm{encode}(\mathrm{decode}(w))\sim w$ up to amalgamation (via HasPushoutSVKEncodeDecode)
@@ 6. Applications
We present $\pi_{1}$ calculations using several methods. Table 1 summarizes the results, distinguishing between direct SVK applications (pushouts with trivial or simple amalgamation), CW-complex constructions (2-cells attached via SVK), and other methods (products, fibrations).
Table 1. Summary of $\pi_{1}$ calculations
Space Fundamental Group Method SVK?
Direct SVK (pushouts with simple amalgamation)
$A\vee B$ $\pi_{1}(A)\ast\pi_{1}(B)$ Pushout over $\mathbf{1}$ Yes
$S^{1}\vee S^{1}$ $\mathbb{Z}\ast\mathbb{Z}$ Wedge of circles Yes
$\bigvee_{n}S^{1}$ $F_{n}$ (free group) $n$-bouquet pushout Yes
$S^{n}$ ($n\geq 2$) 11 Suspension pushout Yes
CW-complex via SVK (2-cell attachment)
$K$ (Klein bottle) $\mathbb{Z}\rtimes\mathbb{Z}$ $S^{1}\vee S^{1}\cup_{aba^{-1}b}D^{2}$ Yes
$\mathbb{RP}^{2}$ $\mathbb{Z}_{2}$ $S^{1}\cup_{a^{2}}D^{2}$ Yes
$\Sigma_{g}$ (genus $g$) Surface group $\bigvee_{2g}S^{1}\cup_{\prod[a_{i},b_{i}]}D^{2}$ Yes
$N_{n}$ (crosscap $n$) $\langle a_{i}\mid\prod a_{i}^{2}=1\rangle$ $\bigvee_{n}S^{1}\cup_{\prod a_{i}^{2}}D^{2}$ Yes
$L(p,q)$ $\mathbb{Z}/p\mathbb{Z}$ Heegaard (tori gluing) Yes
Other methods (not SVK)
$T^{2}$ (torus) $\mathbb{Z}\times\mathbb{Z}$ Product: $\pi_{1}(A\times B)$ No
$\mathbb{CP}^{n}$ 11 Fibration exact sequence No
Remark 6.1 (Scope of Applications).
While this paper focuses on SVK, our formalization includes other fundamental group techniques:
  • Product formula: $\pi_{1}(A\times B)\simeq\pi_{1}(A)\times\pi_{1}(B)$ (used for torus)
  • Fibration sequences: $\pi_{2}(B)\to\pi_{1}(F)\to\pi_{1}(E)\to\pi_{1}(B)$ (used for $\mathbb{CP}^{n}$)
  • Direct encode-decode: For HITs with explicit path constructors (Circle, Klein bottle)
The torus and $\mathbb{CP}^{n}$ are included to show the breadth of the formalization, though they do not use SVK.
@@ 6.1. The Wedge Sum: Free Product of Fundamental Groups
The wedge sum is the pushout $\mathrm{Pushout}(A,B,\mathbf{1})$. Since $\pi_{1}(\mathbf{1})=1$, the amalgamation is trivial, yielding the free product. ∎
@@
1noncomputable def wedgeFundamentalGroupEquiv
2 [HasWedgeSVKDecodeBijective a0 b0] :
3 SimpleEquiv (pi_1(Wedge A B a0 b0, wedgeBase))
4 (FreeProduct (pi_1(A, a0)) (pi_1(B, b0)))
@@ 6.2. The Figure-Eight Space
The figure-eight has a non-abelian fundamental group:
@@
5 intro h; cases h -- constructors are distinct
Remark 6.4 (Non-Commutativity).
The fact that $ab\neq ba$ in $\mathbb{Z}\ast\mathbb{Z}$ demonstrates that $\pi_{1}(S^{1}\vee S^{1})$ is non-abelian. This is a key distinguishing feature from the torus, where $\pi_{1}(T^{2})\cong\mathbb{Z}\times\mathbb{Z}$ is abelian.
6.3. Spheres
Theorem 6.5 (Fundamental Group of $n$-Sphere).
For $n\geq 2$:
$\pi_{1}(S^{n})\simeq 1$
@@ 6.3. Spheres
The $n$-sphere is the (n1)(n-1)-fold iterated suspension: $S^{n}=\Sigma^{n-1}(S^{1})$. By SVK applied to suspension:
$\pi_{1}(\Sigma A)\simeq 1\ast_{\pi_{1}(A)}1=1$
since both “north” and “south” components are points with trivial $\pi_{1}$. ∎
6.4. The Torus (Product Formula, Not SVK)
Theorem 6.6 (Fundamental Group of Torus).
$\pi_{1}(T^{2})\simeq\mathbb{Z}\times\mathbb{Z}$
Remark 6.7 (Torus Construction—Not SVK).
The torus does not use SVK. Instead, we construct $T^{2}:=S^{1}\times S^{1}$ and apply the product formula:
$\pi_{1}(A\times B,(a_{0},b_{0}))\simeq\pi_{1}(A,a_{0})\times\pi_{1}(B,b_{0})$
This requires no additional axioms beyond Circle. The product formula is proved directly from the definition of paths in product types:
$\mathrm{Path}((a_{1},b_{1}),(a_{2},b_{2}))\simeq\mathrm{Path}(a_{1},a_{2})\times\mathrm{Path}(b_{1},b_{2})$
We include the torus here to demonstrate that the formalization supports multiple $\pi_{1}$ calculation techniques, not just SVK.
6.5. The Klein Bottle
Theorem 6.8 (Fundamental Group of Klein Bottle).
$\pi_{1}(K)\simeq\mathbb{Z}\rtimes\mathbb{Z}=\langle a,b\mid aba^{-1}b=1\rangle$
where the semidirect product has $\mathbb{Z}$ acting on $\mathbb{Z}$ by negation.
In Lean:
1structure KleinBottlePresentation where
2 -- Elements are pairs (m, n) where:
3 -- m represents powers of the loop that reverses orientation
4 -- n represents powers of the loop that preserves orientation
5 m : Int
6 n : Int
7
[…]
9-- (m1, n1) * (m2, n2) = (m1 + m2, (-1)^m2 * n1 + n2)
6.6. The Real Projective Plane
Theorem 6.9 (Fundamental Group of Projective Plane).
$\pi_{1}(\mathbb{RP}^{2})\simeq\mathbb{Z}_{2}=\langle a\mid a^{2}=1\rangle$